X's current stock is $10 per share. X will not pay dividends and faces the following conditional prices in a year:

State Probability Price
Good 0.3 $20
Medium 0.4 $10
Low 0.3 $5

In the Good state in a year, Y's conditional price is $5 with an associated 25% return. Y will pay a dividend of $1 with certainty in a year (with Probability=1).

The equally weighted portfolio (50%/50%) that includes Y and X will be riskless if Y's conditional prices in a year are:

A $15 in medium, $10 in low

B $5 in medium, $5 in low

C $15 in medium, $20 in low

D $10 in medium, $15 in low
Answer #1

Option C $15 in medium, $20 in low

Riskless occurs when in all the states we get the same value of the portfolio

As in Good State, the value of the portfolio=value of X+Value of Y=20+5=25

So, in all states value must be 25

In Medium, Value of Y should be =25-10=15
In Low, Value of Y should be =25-5=20